OpenStudy (anonymous):

What is the coefficient of x^n in the expansion of :

OpenStudy (anonymous):

\[(1+x+2x^2+3x^3+...+nx^n)^2\]

OpenStudy (michele_laino):

Please have you tried to use the induction principle?

OpenStudy (irishboy123):

this is (1-x)^(-2) so expand (Binomial) (1-x)^(-4).

OpenStudy (michele_laino):

I got this formula: \[1 + 2{x^2} + 3{x^3} + ... + n{x^n} = n{S_n} - \sum\limits_{n = 1}^{n - 1} {{S_k}} ,\quad with:\;{S_n} = 1 + x + {x^2} + ... + {x^n}\]

ganeshie8 (ganeshie8):

\[\begin{align}(1+x+2x^2+3x^3+...+nx^n)^2 &= \left(1+\sum\limits_{k=1}^{n} kx^{k}\right)^2 \\~\\ &= \left( 1+\sum\limits_{k=1}^{n} x(x^{k})'\right)^2\\~\\ &= \left( 1+x\left[\sum\limits_{k=1}^{n} x^{k}\right]'\right)^2\\~\\ & = \left(1+ x^2\left[\dfrac{x^n-1}{x-1}\right]'\right)^2\\~\\ &= \cdots\end{align}\]

OpenStudy (anonymous):

Thank you everybody ! :)